a document is equally likely to be in any of the 3 box files.a seach of ith the box will discover the document with probability pi. what is the probability that, the document is in the 1st box given that i have searched the first box twice and not found it


1
Expert's answer
2021-07-22T10:16:08-0400

Solution:

There are 3 boxes.

The probability that document will be in any of them"=\dfrac13"

Given, a search of ith box will discover the document with probability pi.

So, a search of 1st box will discover the document with probability = p(1)=p.

Now, the probability that, the document is in the 1st box given that I have searched the first box twice and not found it=P(the document is in the 1st box and I have searched the first box twice and not found it)/P(I have searched the first box twice and not found it)

"=\dfrac{\dfrac13}{(1-p)(1-p)}=\dfrac{1}{3(1-p)^2}"